Operations Report – April 27th

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: A couple sim suit chargers, and soldering iron.

Notes on non-nominal systems: A sim-suit charging cable broke yesterday, due to having cold solder joints. I attempted to repair, but the soldering Iron broke after heating briefly.

Generator (hours run): 9:30 pm to 9:00 am.

Solar— SOC 22 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134192.9

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Yes, 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0.01 Gallons

# Hours the ATVs were Used today: 0.1

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 96 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Eric made progress troubleshooting the beartooth devices by communicating with the manufacturer. The issue has not been resolved yet. Last night we waited too long to turn on the generator, SoC had dropped to 22%. This morning, it was only 26% after the generator running all night. It appears that the generator is not charging the batteries at all. SoC returned to 100% after a few hours of full sunlight. Mission Control (Shannon) is investigating further.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. I changed some settings on radios we use during EVA (whisper function) to make the microphone more sensitive. Recieved new radios, opened one set. They are ready to use on EVAs. Demoted some older EVA radios to science dome, retired one radio from science dome and one from the green hab (both radios have issues).

Summary of Hab operations: Started sorting tools in toolbox, and radios/chargers. Went out and checked water tank, confirmed we are half way through it.

Summary of GreenHab operations: Heavily watered (8+ gallons) in morning at request of mission control.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: New radios retrieved from RAM airlock.

Summary of health and safety issues: During a local EVA to practice medical emergencies, we had one crewmember suffer from heat exhaustion. EVA was cut short, crewmember is feeling better.

Questions, concerns and requests to Mission Support: We would like some small cardboard boxes for organizing the HAB. They would also be good for moving material and supplies during EVAs (such as to/from RAM).
We would like some hang-tags or white cloth (gaffing?) tape to mark things as we organize the HAB.
Also, we would like some small paintbrushes (for detail work on sign), wet-wipes, and paper towels.

Copyright © The Mars Society. All rights reserved. | Main Site